OpenStudy (anonymous):

cos3x=cos^3(x)-3sin^2(x)Cos(x)

OpenStudy (anonymous):

could you use the Equation editor instead, it's not very clear above with the powers?

OpenStudy (anonymous):

\[\cos 3x=\cos ^{3}x-3\sin ^{2}x \cos x\]

OpenStudy (anonymous):

what's required? simplify or find x?

OpenStudy (anonymous):

prove idenity

OpenStudy (anonymous):

cos(3x) = cos(2x + x) = cos(2x)cos(x) - sin(2x)sin(x) = [cos(x)^2 - sin(x)^2] * cos(x) - 2sin(x)cos(x)sin(x) = cos(x)^3 - sin(x)^2 * cos(x) - 2sin(x)^2 * cos(x) = cos(x)^3 - 3sin(x)^2 * cos(x)